# GLDesigner: Leveraging Multi-Modal LLMs as Designer for Enhanced Aesthetic Text Glyph Layouts

### v1

@@ Abstract
Text logo design heavily relies on the creativity and expertise of professional designers, in which arranging element layouts is one of the most important procedures. However, few attention has been paid to this specific task which needs to take precise textural details and user constraints into consideration, but only on the broader tasks such as document/poster layout generation. In this paper, we propose a VLM-based framework that generates content-aware text logo layouts by integrating multi-modal inputs with user constraints, supporting a more flexible and stable layout design in real-world applications. We introduce two model techniques to reduce the computation for processing multiple glyph images simultaneously, while does not face performance degradation. To support instruction-tuning of out model, we construct two extensive text logo datasets, which are 5x more larger than the existing public dataset. Except for the geometric annotations (e.g. text masks and character recognition), we also compliment with comprehensive layout descriptions in natural language format, for more effective training to have reasoning ability when dealing with complex layouts and custom user constraints. Experimental studies demonstrate the effectiveness of our proposed model and datasets, when comparing with previous methods in various benchmarks to evaluate geometric aesthetics and human preferences. The code and datasets will be publicly available.
1 Introduction
@@
Automatic layout design in creative AI is gaining attention for boosting efficiency and supporting designers. Traditional rule-based methods often produce plain, unappealing layouts and rely on professional designers for aesthetic results, making the process labor-intensive. Particularly, text logo layout designing is challenging and less explored than graphic designs because it requires careful consideration of font style, size, texture and keyword semantics, to ensure the final design is both visually appealing and collision-free.
@@
Figure 1: Illustrated are examples of text logos from our datasets, each glyph uniquely colored. We aim to enhance previous datasets by introducing more complex, compact layouts (top) and extended text sequences (bottom).
@@
To overcome these challenges, recent approaches leverage generative models such as Generative Adversarial Networks (GANs) [10] and Diffusion Models [35] to synthesize graphic layouts by learning from examples of human-designed data. The previous method [38] utilized GAN-based models to address this issue but remains limited in handling (1) long text sequences, (2) user-defined layout constraints, and (3) data variability and generality, making it less suitable for real-world design applications. Additionally, the method encodes character semantics using pretrained semantic word embedding models [26], which are less expressive compared to contemporary large language models (LLMs) that are trained on trillions of text tokens. Meanwhile, an increasing number of general layout generation methods [9, 36, 28, 41] have begun exploring multi-modal, content-aware conditions (e.g., background images, document contexts). However, these approaches cannot be directly applied to the text logo layout task, as it requires managing multiple glyphs and incorporating more fine-grained details to ensure both aesthetic appeal and readability.
@@
Acknowledging these limitations, we present a unified VLM-based framework for text logo layout generation, leveraging their inherent ability to handle multi-modal inputs of glyph images and textual semantics. We employ pre-trained VLMs [29] with minor modifications, including early feature extraction to enhance sensitivity to fine-grained textural details and prevent collisions between glyphs. To improve computational efficiency, we apply adaptive average pooling, reducing the number of visual glyph tokens by a factor of $6^{2}$, which accelerates both training and inference without adding extra parameters or compromising performance. The final layout outputs are represented as pure JSON texts, ensuring a standardized format that facilitates easy and stable post-rendering while enhancing training stability. Additionally, we fine-tune the LLM with instruction-following techniques for the text logo layout task, enabling it to create aesthetically pleasing layouts by analyzing glyph styles, semantic meanings, and optionally incorporating user constraints.
@@
We construct two large-scale text logo datasets featuring pixel-level annotations, detailed layout descriptions, and user constraints, comprising over 17k samples, which is six times larger than existing public datasets [38]. These datasets include synthesized logos rendered by pre-defined rules from human designers and general text logos sourced from the internet, which are subsequently subjected to post-processing and annotation, as illustrated in Fig. 1. Extensive experiments demonstrate that leveraging both datasets significantly reduces glyph collisions and enhances layout creativity and diversity. Our approach outperforms all existing methods on qualitative benchmarks and human preference evaluations while also satisfying additional layout constraints, enabling more controllable generation in real-world applications.
@@
    We introduce a VLM-based framework that generates content-aware text logo layouts by integrating multi-modal inputs and optionally supports user constraints through natural language instructions, making it well-suited for real-world design scenarios.
  2. 2.
@@
    We construct two extensive text logo datasets, comprising synthesized and general samples respectively, and generate instruction-following data with fine-grained layout descriptions, significantly surpassing previous efforts in both scale and detail.
  3. 3.
@@
    Our method outperforms existing approaches across multiple benchmarks, and effectively handles long text sequences and user-defined layout requirements. By combining with glyph transformation and textural generation, our model showcases its versatility in real-world applications.
2 Related Work
2.1 Automatic Layout Generation
@@
Graphic Layout Generation encompasses neural networks to automatically arrange elements like images, text, and backgrounds into appealing designs. Early models such as LayoutGAN [24] and LayoutVAE [20] focused on document and poster arrangement without specific content considerations, utilizing generative models [10, 22]. Recent approaches with auto-regressive [2, 11, 23] and diffusion models [18, 44, 17] enhance layout realism and diversity but often lack the content specificity needed for real-world applications. In contrast, content-aware layout generation has grown to address more complex design scenarios, especially in commercial poster design with non-empty background images, guided by saliency maps [14, 46]. Conditional models integrate visual and textual encoders to predict layouts based on specific conditions [4, 43], while other methods employ constrained optimization [21] and intermediate representations [19, 27] to adhere to user-specified constraints.
@@
Text Logo Layout Generation is a specialized area within graphic design, focusing exclusively on arranging pixel-level text glyphs, which requires careful consideration of glyph strokes and reading order to ensure both aesthetic appeal and readability. Currently, the only existing method [38] utilizes a GAN-based model to integrate both linguistic and visual information from input texts and glyph images to predict optimal layouts. While demonstrating promising performance, it still struggles with designing layouts for long text sequences, accommodating user-defined constraints, and offering diverse layout designs due to the limited amount of training data.
@@
Figure 2: Pipeline Overview of the Layout Generation Method. For clarity and simplicity, vision-language projection layers and detailed prompt templates are not depicted. Utilizing Early Feature Fusion and Adaptive Average Pooling techniques, as detailed in Section 3.2, ensures precise recognition of the glyph image’s textural details with a reduced number of glyph visual tokens. The JSON-structured output guarantees a consistent and reliable rendering of the final text logo.
2.2 Multi-modal Large Language Models
@@
While layouts, which can be encoded in coding formats such as HTML or JSON, are ideally suited to be processed by pre-trained Large Language Models (LLMs), as domain-specific data can be utilized to strength the code generation ability. Following similar idea, several works use the same method to finetune VLLMs to have better performance on document layout understanding [31, 16]. LayoutNUWA [36] views layout generation as a code generation task and LayoutPrompter [28] leverages the training-free RAG pipeline to strengthen the in-context learning ability. Recently, PosterLLaVA [41] employs structured text and visual instruction tuning to generate layouts under specific visual and textual constraints for poster designs. Meanwhile, some contemporaneous works [6, 5, 9] adopt LLMs as middle-ware for designing and direct layout-to-image generation through diffusion models.
@@ 3.1 Preliminaries
A text logo can be represented as a collection of $N$ glyph elements, denoted by $\{(g_{i},b_{i})\}^{N}_{i=0}$. Here, each $g_{i}\in\mathbb{R}^{H_{i}\times W_{i}}$ represents the raw visual component of the glyph image. The corresponding position $b_{i}=(x_{i},y_{i},x_{i}^{\prime},y_{i}^{\prime})$ specifies the glyph’s location using a bounding box defined by its top-left $(x_{i},y_{i})$ and bottom-right $(x_{i}^{\prime},y_{i}^{\prime})$ coordinates. These coordinates are normalized relative to the background’s overall width and height, ensuring consistent placement across different resolutions.
@@
In the task of generating text logo layouts, the inputs include each glyph’s character text $T_{i}$, its corresponding glyph image $g_{i}$, and optionally, user-defined constraints $c$. The objective is to predict the positional bounding box $b_{i}$ for each glyph based on these inputs.
@@
Table 1: Prompt templates for applying instruction fine-tuning on the content-aware text logo generation task. The placeholder $T_{i}$ and <glyphi> will be replaced with character texts and glyph visual features during training and inference. Note this is one example of the prompt template, and we have an extensive prompt lists offering unique descriptions of layout-related questions, which will be randomly selected during training to enhance generalization and adaptability.
3.2 Model Architecture
We provide a comprehensive overview of our layout generation pipeline in Fig. 2, with detailed explanations elaborated in the subsequent subsections. The architecture of our model is built upon LLaVA [29], which facilitates the unified processing of glyph images and textual content. Specifically, we utilize ViT-L [33] as the glyph encoder and LLaMA [37] as the LLM. To further enhance the model’s performance, we introduce two innovative techniques: Early Feature Fusion and Adaptive Average Pooling. The first technique enhances the model’s ability to capture and interpret intricate glyph textural details, such as stroke patterns and appearance characteristics. Concurrently, The latter boosts computational efficiency when processing multiple glyph images by reducing the number of glyph visual tokens. These enhancements collectively empower our model to more effectively understand and generate precise layout configurations.
Early Feature Fusion. The default configuration of LLaVA [29] utilizes the visual features from CLIP’s [33] second-to-last layer for vision-language feature projection. While these features focus more on capturing the global properties and semantics of glyph images, but tend to overlook the finer textural details present in the earlier layers. Therefore, we choose to fuse the early visual features into the original features, and experimentally and symmetrically choose the second CLIP layers. We also add a trainable two-layer MLP as the projection layer, which is zero-initialized for training stabilization.
Adaptive Average Pooling. Another limitation of the original model is the linear increase in the number of visual tokens when processing multiple glyph images simultaneously. The number of processed tokens can easily exceed model’s maximum sequence length, leading to decreased training and inference efficiency. To address this issue, we utilize adaptive average pooling to compress the number of visual tokens without introducing additional parameters beyond the existing trainable modules [25, 3]. As demonstrated in [40, 42], this is a parameter-free technique that preserves the spatial locality while retaining essential visual attributes. Moreover, our experimental results confirm that this approach significantly reduces computational overhead during both training and inference phases, all without compromising the quality of layout generation.
3.3 Layout Structure
We introduce tailored prompt templates and a structured textual layout arrangement in JSON format, crafted specifically for the instruction-following fine-tuning stage. Details of the template are provided in Tab. 1.
To effectively guide the model, we begin with a detailed task description in the instruction prompt, which serves to clearly indicate and specify both the task content and the desired output format. For handling glyph images alongside their textual semantics, each glyph is represented with unified tokens as $T_{i}$: <glyphi>, where $T_{i}$ and <glyphi> represent the the $i$-th text character and the visual features of the glyph, as extracted in Sec. 3.2, respectively. Finally, we append an initial empty JSON structure that retains only the essential keys. This minimalist JSON setup helps streamline the training process by providing a clear and uncluttered framework for the model to learn within.
To construct the ground-truth response answers, we enhance the JSON structure by incorporating a comprehensive global layout description alongside a “detail” property. The “detail” property specifies the detailed description for each individual glyph, leveraging the chain-of-thought methodology [39]. This approach facilitates the management of complex and lengthy sequence layouts through a systematic, step-by-step reasoning process. This enhancement not only boosts the accuracy of intricate layout generation but also ensures greater flexibility and responsiveness to a diverse range of user requirements.
4 Datasets
Dataset Train Test Glyphs/img Total Glyphs
TextLogo3K [38] 3170 300 4.69 16274
SynTextLogo 10000 1000 4.36 47963
GenTextLogo 6586 482 5.24 37036
Table 2: Dataset statistics of existing dataset and our datasets. Our proposed datasets boast a significantly larger scale than the existing one, featuring a higher average and total number of glyphs, specifically designed to address complex and lengthy sequence scenarios that [38] struggles to handle.
Figure 3: Textual Layout Description complemented by our data processing pipeline in Sec. 4.3. We provide detailed textual descriptions for the overall text logo (top) and each text glyphs (bottom), with geometric relationships between glyphs.

### v2

@@ Abstract.
Text logo design heavily relies on the creativity and expertise of professional designers, in which arranging element layouts is one of the most important procedures. However, this specific task has received limited attention, often overshadowed by broader layout generation tasks such as document or poster design. In this paper, we propose a Vision-Language Model (VLM)-based framework that generates content-aware text logo layouts by integrating multi-modal inputs with user-defined constraints, enabling more flexible and robust layout generation for real-world applications. We introduce two model techniques that reduce the computational cost for processing multiple glyph images simultaneously, without compromising performance. To support instruction tuning of our model, we construct two extensive text logo datasets that are five times larger than existing public datasets. In addition to geometric annotations (e.g., text masks and character recognition), our datasets include detailed layout descriptions in natural language, enabling the model to reason more effectively in handling complex designs and custom user inputs. Experimental results demonstrate the effectiveness of our proposed framework and datasets, outperforming existing methods on various benchmarks that assess geometric aesthetics and human preferences.
Text Glyph Layout, Layout Generation, Vision-Language Models
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3755289ccs: Computing methodologies Natural language generationccs: Applied computing Media artsccs: Human-centered computing User centered design
1. Introduction
@@
Automatic layout design in creative AI is gaining increasing attention for its potential to enhance efficiency and assist designers in the creative process. Traditional rule-based methods often result in plain and aesthetically unappealing layouts, heavily relying on professional designers to achieve visually pleasing outcomes—making the process time-consuming and labor-intensive. In particular, text logo layout design poses unique challenges and remains underexplored compared to general graphic layout tasks. It requires careful consideration of font style, size, texture, and the semantic importance of keywords to ensure that the final design is both visually compelling and free of glyph collisions.
@@
Figure 1. Illustrated are examples of text logos from our datasets, each glyph uniquely colored. We aim to enhance previous datasets by introducing more complex, compact layouts (top) and extended text sequences (bottom).
@@
To overcome these challenges, recent approaches leverage generative models such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) and Diffusion Models (Sohl-Dickstein et al., 2015) to synthesize graphic layouts by learning from examples of human-designed data. The previous method (Wang et al., 2022) utilized GAN-based models to address this issue but remains limited in handling (1) long text sequences, (2) user-defined layout constraints, and (3) data variability and generality, making it less suitable for real-world design applications. Additionally, the method encodes character semantics using pretrained semantic word embedding models (Li et al., 2018), which are less expressive compared to contemporary large language models (LLMs) that are trained on trillions of text tokens. Meanwhile, an increasing number of general layout generation methods (Feng et al., 2024; Tang et al., 2023; Lin et al., 2023b; Yang et al., 2024) have begun exploring multi-modal, content-aware conditions (e.g., background images, document contexts). However, these approaches cannot be directly applied to the text logo layout task, as it requires managing multiple glyphs and incorporating more fine-grained details to ensure both aesthetic appeal and readability.
@@
Acknowledging these limitations, we propose a unified VLM-based framework for text logo layout generation, leveraging the inherent capability of vision-language models to process multi-modal inputs, including glyph images and textual semantics. Our approach builds upon pre-trained VLMs (Liu et al., 2023) with minimal architectural modifications—specifically, early-stage feature extraction to improve sensitivity to fine-grained textural details and mitigate glyph collisions. To enhance computational efficiency, we apply adaptive average pooling, reducing the number of visual glyph tokens by a factor of $6^{2}$. This significantly accelerates both training and inference without introducing additional parameters or degrading performance. The final layout outputs are represented as pure JSON texts, ensuring a standardized format that facilitates easy and stable post-rendering while enhancing training stability. Furthermore, we fine-tune the language component using instruction-following techniques tailored to the text logo layout task, enabling the model to generate aesthetically pleasing designs by reasoning over glyph styles, semantic content, and optional user-defined constraints.
@@
We construct two large-scale text logo datasets containing pixel-level annotations, detailed layout descriptions, and user-defined constraints, totaling over 17,000 samples — five times larger than the currently available public dataset (Wang et al., 2022). The datasets comprise synthesized logos generated using rule-based methods derived from expert designers, as well as real-world text logos collected from the internet, followed by systematic post-processing and annotation, as illustrated in Figure 1. Extensive experiments demonstrate that leveraging both datasets significantly reduces glyph collisions and enhances layout creativity and diversity. Our proposed approach outperforms all existing methods across qualitative benchmarks and human preference studies, while also accommodating additional layout constraints—enabling more controllable and practical text logo generation in real-world scenarios.
@@
    We propose a VLM-based framework for content-aware text logo layout generation, which integrates multi-modal inputs and optionally supports user-defined constraints via natural language instructions, enabling greater flexibility in real-world design scenarios.
  2. (2)
@@
    We construct two large-scale text logo datasets, comprising both synthesized and real-world samples, with fine-grained layout annotations and instruction-following data. These datasets significantly surpass existing ones in both scale and descriptive richness.
  3. (3)
@@
    Our method outperforms existing approaches across multiple benchmarks, effectively handling long text sequences and accommodating user-defined layout requirements. Furthermore, by incorporating glyph transformation and texture generation, our model demonstrates strong versatility in practical applications.
2. Related Work
2.1. Automatic Layout Generation
@@
Graphic Layout Generation encompasses neural networks to automatically arrange elements such as images, text, and backgrounds into visually appealing designs. Early models like LayoutGAN (Li et al., 2019) and LayoutVAE (Jyothi et al., 2019) focused on document and poster layouts without considering specific content, relying on generative frameworks such as GANs (Goodfellow et al., 2014) and VAEs (Kingma, 2013). Recent advances using auto-regressive models (Arroyo et al., 2021; Gupta et al., 2021; Kong et al., 2022) and diffusion-based approaches (Inoue et al., 2023; Zhang et al., 2023a; Hui et al., 2023) have significantly improved layout realism and diversity. However, these models often lack the content-awareness required for practical applications. In response, content-aware layout generation has emerged to tackle more complex design scenarios, particularly in commercial poster design with non-empty backgrounds, often incorporating saliency maps for layout guidance (Hsu et al., 2023; Zhou et al., 2022). Conditional models further enhance layout prediction by integrating visual and textual encoders (Cao et al., 2022; Yu et al., 2022), while other approaches utilize constrained optimization (Kikuchi et al., 2021) and intermediate representations (Jiang et al., 2023; Lin et al., 2023a) to better conform to user-defined constraints.
@@
Text Logo Layout Generation is a specialized domain within graphic design that focuses on the arrangement of pixel-level text glyphs. This task demands meticulous attention to glyph stroke details and reading order to achieve both aesthetic harmony and legibility. Currently, the only existing method (Wang et al., 2022) utilizes a GAN-based model to integrate both linguistic and visual information from input texts and glyph images to predict optimal layouts. While this method shows promising results, it faces notable limitations in handling long text sequences, supporting user-defined constraints, and producing diverse layout variations—largely due to the scarcity of high-quality training data.
@@
Figure 2. Pipeline Overview of the Text Logo Layout Generation Method. For clarity and simplicity, vision-language projection layers and detailed prompt templates are omitted. By incorporating Early Feature Fusion and Adaptive Average Pooling, as described in Section 3.2, the model ensures precise recognition of the glyph image’s textural details while reducing the number of visual tokens. The JSON-structured output guarantees a consistent and reliable rendering of the final text logo.
2.2. Multi-modal Large Language Models
@@
While layouts, which can be encoded in coding formats such as HTML or JSON, are ideally suited to be processed by pre-trained Large Language Models (LLMs), as domain-specific data can be utilized to strength the code generation ability. Following similar idea, several works use the same method to finetune VLLMs to have better performance on document layout understanding (Luo et al., 2024; Huang et al., 2022). LayoutNUWA (Tang et al., 2023) views layout generation as a code generation task and LayoutPrompter (Lin et al., 2023b) leverages the training-free RAG pipeline to strengthen the in-context learning ability. Recently, PosterLLaVA (Yang et al., 2024) employs structured text and visual instruction tuning to generate layouts under specific visual and textual constraints for poster designs. Meanwhile, some contemporaneous works (Chen et al., 2024, 2023a; Feng et al., 2024) adopt LLMs as middle-ware for designing and direct layout-to-image generation through diffusion models.
@@ 3.1. Preliminaries
A text logo can be represented as a collection of $N$ glyph elements, denoted by $\{(g_{i},b_{i})\}^{N}_{i=0}$. Each $g_{i}\in\mathbb{R}^{H_{i}\times W_{i}}$ represents the raw visual component of the $i$-th glyph in image form, where $H_{i}$ and $W_{i}$ denote its height and width, respectively. The corresponding bounding box $b_{i}=(x_{i},y_{i},x_{i}^{\prime},y_{i}^{\prime})$ defines the glyph’s position, with $(x_{i},y_{i})$ and $(x_{i}^{\prime},y_{i}^{\prime})$ representing the top-left and bottom-right coordinates of the box. These coordinates are normalized with respect to the overall width and height of the canvas, ensuring resolution-independent and consistent placement across varying background sizes.
@@
In the task of generating text logo layouts, the inputs consist of each glyph’s character text $T_{i}$, its corresponding glyph image $g_{i}$, and optionally, user-defined constraints $c$. The goal is to predict the positional bounding box $b_{i}$ for each glyph, based on these inputs, such that the resulting layout is visually coherent, semantically meaningful, and adheres to any specified constraints.
3.2. Model Architecture
We present an overview of our layout generation pipeline in Figure 2, with detailed explanations provided in the following subsections. Our model architecture is built upon LLaVA (Liu et al., 2023), which enables unified processing of glyph images and textual inputs. Specifically, we employ ViT-L (Radford et al., 2021) as the glyph encoder and LLaMA (Touvron et al., 2023) as the language model. To further enhance the model’s performance, we introduce two key techniques: Early Feature Fusion and Adaptive Average Pooling. Early Feature Fusion improves the model’s ability to capture and interpret fine-grained glyph details, including stroke patterns and visual characteristics. Adaptive Average Pooling, on the other hand, reduces the number of visual tokens by downsampling glyph image representations, significantly improving computational efficiency when handling multiple glyphs. Together, these enhancements enable our model to more effectively understand visual-textual inputs and generate precise, aesthetically pleasing layout configurations.
3.2.1. Early Feature Fusion.
The default configuration of LLaVA (Liu et al., 2023) utilizes visual features from CLIP’s (Radford et al., 2021) second-to-last layer for vision-language feature projection. While these features capture the global properties and semantics of glyph images effectively, they tend to overlook finer textural details present in earlier layers. To address this, we propose fusing early visual features into the original feature set. Based on experimental results, we select the second-to-last layer of CLIP as the point of fusion. Additionally, we introduce a trainable two-layer multi-layer perceptron (MLP) as a projection layer, initialized to zero to ensure training stability.
3.2.2. Adaptive Average Pooling.
Another limitation of the original model is the linear increase in the number of visual tokens when processing multiple glyph images simultaneously. As the number of glyph images grows, the total number of tokens can quickly exceed the model’s maximum sequence length, which in turn leads to a reduction in both training and inference efficiency.
To address this issue, we employ adaptive average pooling to reduce the number of visual tokens, without introducing any additional parameters beyond the existing trainable modules (Li et al., 2023; Bai et al., 2023). As demonstrated in (Xu et al., 2024; Yao et al., 2024), adaptive average pooling is a parameter-free technique that effectively preserves spatial locality while retaining key visual features. Our experimental results further confirm that this approach significantly reduces computational overhead during both the training and inference phases, all while maintaining the quality of layout generation.
3.3. Layout Structure
We introduce customized prompt templates and a structured textual layout representation in JSON format, specifically designed for the instruction-following fine-tuning stage. The details of the template are provided in Table 2.
To effectively guide the model, we begin with a detailed task description in the instruction prompt, which serves to clearly indicate and specify both the task content and the desired output format. For handling glyph images alongside their textual semantics, each glyph is represented by unified tokens as $T_{i}$: <glyphi>, where $T_{i}$ represents the $i$-th text character and <glyphi> corresponds to the visual features of the glyph, as extracted in Section 3.2.1. Additionally, we append an initial empty JSON structure that retains only the essential keys. This minimalist JSON setup facilitates the training process by providing a clear and straightforward framework for the model to learn from.
To construct the ground-truth response answers, we enhance the JSON structure by incorporating a comprehensive global layout description alongside a ”detail” property. The ”detail” property specifies the detailed description for each individual glyph, leveraging the chain-of-thought methodology (Wei et al., 2022). This approach facilitates the management of complex and lengthy sequence layouts through a systematic, step-by-step reasoning process. This enhancement not only boosts the accuracy of intricate layout generation but also ensures greater flexibility and responsiveness to a diverse range of user requirements.
Dataset Train Test Glyphs/img Total Glyphs
TextLogo3K (Wang et al., 2022) 3170 300 4.69 16274
SynTextLogo 10000 1000 4.36 47963
GenTextLogo 6586 482 5.24 37036
Table 1. Dataset statistics of existing dataset and our datasets.
@@ 3.3. Layout Structure
Table 2. Prompt templates for applying instruction fine-tuning on the content-aware text logo generation task. The placeholder $T_{i}$ and <glyphi> will be replaced with character texts and glyph visual features during training and inference.
4. Datasets